Help please
25 of 40 (0 complete) Apharmaceutical company clams that a dug s effective in treating 90% of people han g severe m gane headaches What is the probability that exactly 5 people will have relief from ther migraine O A 0354 O B. 0.50 The digsahnsted to 6 erts that have sevn mga e C. 0.059 O D. 0 000054 What is the mean of the binomial distribution OA 54 B. 56 OC 054 D. 55 the standard deviation ot the bromad eistrbtion? A 0.29 B 0.73 C. 054 wat Chck to select your ansover
Answer #1

Question 1:

n = 6, p = 0.9 (because 90%), r = 5

so, probability that 5 out of 6 patients will get relief from the migraine,

pr(1-P)-r = 5 0.93(1 - 0.9)6-50.35-4

So, option A is correct.

Question 2:

Mean of binomial distribution = np = 6 * 0.9 = 5.4. So option A is correct.

Question 3:

Standard deviation of binomial distribution = np(1-p) = 6 * 0.9 * 0.1 = 0.54. So option C is correct.

** If the answer does not match then please comment.